EX 18-16 Cost of production report The debits to Work in Process--Roasting Department for St. Arbucks Coffee Company for Julyplease answer a to e
Answer #1
St. Arbucks Coffee company
FIFO Method:
Equivalent Units
DM Conversion costs
Whole units % Units % Units
Opening WIP 600
Started 23000
Total Units to account for 23600
Opening WIP 600 0 0 80 480
Completed and tfd. 22000 100 22000 100 22000
Closing WIP 1000 100 1000 42 420
Total Units to account for ( EUP) 23600 23000 22900 Answer 1
1 -Cost Information
Costs to account for: Materials Conversion Total
Op. WIP 2280 138 2418
Added during the period 82800 27480 110280
Total costs to account for 85080 27618 112698
Cost per EUP :-
Cost to be accounted for 82800 27480 110280
EUP 23000 22900
Cost per EUP 3.6 1.2 4.8 Answer 2
2 - Costs Accounted for: Materials Conversion Total
Completed and tfd.
Opening WIP 2280 138 2418
Completed and tfd. 22000 * 3.6 79200 22480 * 1.2 26976 106176
Completed and tfd. 81480 27114 108594
Cl. WIP 1000 * 3.6 3600 420 *1.2 504 4104
Total costs accounted for 85080 27618 112698
Cost reconciliation schedule
Costs Accounted for:
Transferred out 108594 Answer 3
closing WIP
Materials 3600
Conversion costs 504 4104 Answer 4
Total costs 112698
Answer 5
Opening WIP contains Material cost of $ 2280 and 600 units were completed so mat. Cost p.u = 2280 / 600 = 3.8, there is a decrease this time
( 3.8 - 3.6 ) i.e $ 0.2
Conversion cost per unit = 138 / 120 = 1.15, there is an increase this time ( 1.2 - 1.15 ) i.e $ 0.05
Since the format was not available, the question has been solved through this format.